A stock price is currently $100. Over each of the next two 6-month periods it is expected to go up by 10% or down by 10%. The risk-free interest rate is 8% per annum with continuous compounding. What is the value of a 1-year European call option with a strike price of $100?
Answer #1

Step 2 Step 5 Scenario price-Spot price*(U)^2=100*1.1^2=121 Binomial Tree Pricing Payoff HH-Max(Scenario price-Strike price,0